# Approximate Bayesian Computation sequential Monte Carlo via random forests

### v1

@@ Abstract
Approximate Bayesian Computation (ABC) is a popular inference method when likelihoods are hard to come by. Practical bottlenecks of ABC applications include selecting statistics that summarize the data without losing too much information or introducing uncertainty, and choosing distance functions and tolerance thresholds that balance accuracy and computational efficiency. Recent studies have shown that ABC methods using random forest (RF) methodology perform well while circumventing many of ABC’s drawbacks. However, RF construction is computationally expensive for large numbers of trees and model simulations, and there can be high uncertainty in the posterior if the prior distribution is uninformative. Here we adapt distributional random forests to the ABC setting, and introduce Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). This updates the prior distribution iteratively to focus on the most likely regions in the parameter space. We show that ABC-SMC-(D)RF can accurately infer posterior distributions for a wide range of deterministic and stochastic models in different scientific areas.
@@ 1 Introduction
Mathematical modeling has played an important role in studying scientific phenomena. Its practical applications often depend on accurately extracting model parameters $\theta$ from experimentally observed data, $y^{obs}$. In the Bayesian framework, this entails inferring the posterior distribution
$\pi(\theta\mid y^{obs})\propto f(y^{obs}\mid\theta)\cdot\pi(\theta)$
@@
of the parameters from the data, exploiting the likelihood $f(y^{obs}\mid\theta)$ of such observations under the model with given parameter values. However, the likelihood function is often intractable to derive theoretically, difficult to compute numerically, or is too complex to optimize directly. Approximate Bayesian Computation (ABC) was proposed as an alternative method to approximate the posterior distributions in such scenarios (Tavaré et al., 1997; Fu and Li, 1997; Pritchard et al., 1999; Beaumont et al., 2002). Using statistics to summarize model simulations, ABC seeks the parameters that result in minimal distance between the statistics and those of the observed data. Different variations have been proposed to improve the performance of the original ABC method, such as ABC Monte Carlo Markov Chain (Marjoram et al., 2003) and ABC Sequential Monte Carlo (Toni et al., 2008), and they have collectively found remarkable success in applications across different scientific areas (Sisson et al., 2018).
@@
The accuracy and efficiency of ABC in practice depend on several factors. First, it requires a distance function to compare observed and simulated data. The metric is typically weighted so that different statistics contribute equally (Prangle, 2017). However, an equally important consideration is the relevance of each statistic, defined as the amount of information that it carries toward identifying the underlying parameters. Optimizing these two criteria can be challenging (Jung and Marjoram, 2011). Second, ABC requires a tolerance threshold to decide whether proposed parameters are accepted or rejected. The tolerance $\epsilon$ poses a trade-off between the computational efficiency, which increases with higher $\epsilon$, and accuracy, which improves as $\epsilon$ decreases, of the empirical posterior distributions. Choosing $\epsilon$ depends on the specific model, statistic choices and distance function, and it may require intuition or experimentation (Jung and Marjoram, 2011). Finally, and most importantly, ABC’s results depend on the choice of summary statistics. The approximated posterior distribution is only guaranteed to converge to the likelihood-based distribution for small $\epsilon$ if the statistics are sufficient. For complex models, it is not always possible or realistic to find low-dimensional sufficient statistics (Sisson et al., 2018, Chapter 5). This can sometimes be remedied by inclusion of many distinct statistics. However, the posterior estimation may become distorted if the selected statistics are noisy or uninformative. These factors in combination mean that optimizing the performance of ABC in practice can be challenging and require extensive experimentation.
@@
We review some traditional ABC methods in Appendix A, and use them in comparisons with ABC-SMC-(D)RF described in the following sections. Recent ABC implementations using random forests are reviewed in Section 2, with illustrative examples. In Section 4, we describe ABC-SMC-(D)RF’s methodology. We then demonstrate its performance, compared both to traditional ABC algorithms and previous RF-based methods, in a wide range of deterministic and stochastic models (Section 5). We conclude with a discussion.
@@ 2 Random forests in the context of ABC
Although simple to implement, the ABC methods described in Appendix A depend heavily on the choices of hyperparameters, including the summary statistic function $S(y)$, the distance function $d(s,s^{\prime})$ and the tolerance threshold $\epsilon$ (Sisson et al., 2018). Furthermore, the choices of Markov kernels to propose new particles in ABC-MCMC and ABC-SMC also affect the approximated posterior distributions. As random forests (RF) gain in popularity as a powerful non-parametric regression technique (Breiman, 2001; Segal, 2004; Rigatti, 2017; Desai and Ouarda, 2021), its applications in the context of ABC are also becoming more prominent. RF-based methods eliminate the need for choice of metric, tolerance level and the perturbation kernels, making them an attractive alternative to the previous ABC implementations. Importantly, the RF method is relatively robust to noise (Marin et al., 2018), in that $S(y)$ can include many poorly informative statistics without significant impact on the results.
@@ 2.1 One-dimensional ABC random forests
Raynal et al. introduced ABC random forest (ABC-RF) to approximate the posterior distribution for ABC inference problems with one parameter (Raynal et al., 2019), based on the random forest formulation (Algorithm 1) developed by Breiman (2001). The root node for each tree $T_{t}$ is the reference table, consisting of parameters drawn from the prior distribution and simulated statistics. Each node $R_{*}$ in the decision tree is split into two nodes $R_{*1}$ and $R_{*2}$ by selecting one statistic $k_{*}\in\{1,\dots,|S|\}$, where $|S|$ is the total number of summary statistics, and a threshold $s_{*}$ such that the $L^{2}$-loss between the divided sets
@@
is minimized, where $|R_{*}|$ is the number of particles in $R_{*}$, and $\bar{\theta}_{1}$, $\bar{\theta}_{2}$ are the parameter means in $R_{*1}$ and $R_{*2}$ respectively. The tree is finished when each leaf either consists of less than $N_{min}$ particles, or all of the leaf’s particles have the same statistics. To guarantee forest diversity, the root node of each tree (Step 5 in Algorithm 1) is bootstrapped from the reference table, and the statistics considered for splitting each node (Step 9 in Algorithm 1) is limited to a randomly selected subset of size $n_{try}$ among the available statistics $S$.
@@ 6      Tree $T_{t}$ is rooted in node $R_{0}$
7      while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{min}$ do
8             Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{min}$
9            Statistic candidates for splitting $R_{*}\leftarrow$ sample of size $n_{try}$ among $S$
@@
Given an observation $y^{obs}$, predicting its parameter $\theta$ based on tree $T_{t}$ involves following from the root node with the statistics $s^{obs}$ and comparing $s^{obs}_{k_{*}}$ with $s_{*}$ at each node $R_{*}$ to select the next node, until reaching a leaf $L_{t}(s^{obs})$. The prediction for $\theta$ based on $T_{t}$ is then the average parameter among particles in $L_{t}(s^{obs})$:
$\frac{\sum_{i=1}^{N}\theta^{(i)}\cdot n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
where $n_{t}^{(i)}$ is the number of times $\theta^{(i)}$ is duplicated in the bootstrapped sample $R_{0}$, and $\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}$ is the indicator for whether $s^{(i)}$ falls into the same leaf as $s^{obs}$. The prediction for $\theta$ based on the whole forest is the average of predictions based on each tree:
$\sum_{i=1}^{N}\theta^{(i)}\cdot\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
Raynal et al. (2019) argued that this weighted estimate implies that the density of particles $\theta^{(i)}$ with corresponding weights $w^{(i)}$ (Algorithm 2) forms the approximation $\pi_{ABC-RF}\left(\theta\middle|s^{obs}\right)$ for the posterior distribution.
@@
2       Follow the tree $T_{t}$ from root node with $s^{obs}$ until locating leaf $L_{t}(s^{obs})$
@@
4       Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
@@
Marin et al. (2022) developed an R package abcrf based on the ABC-RF method. For multivariate problems, ABC-RF is usually applied for each parameter marginally (Raynal et al., 2019). The authors also provided suggested values for RF hyperparameters; by default, $N_{min}=5$ and $n_{try}=|S|/3$. The most important hyperparameter is the count of trees $B$, for which they recommend analyzing whether the out-of-bag mean squared error stabilizes around the selected $B$ (Raynal et al., 2019; Pudlo et al., 2016).
2.2 An example from population genetics
We examine the performance of ABC-RF with a classic example from population genetics. We consider a sample of $n=1000$ DNA sequences collected from a stationary population that evolved according to the neutral coalescent model of Kingman (1982) with the infinitely-many-sites mutation model (ISM) (Kimura, 1969) with mutation rate $\theta$. Two statistics that are often used to summarize the mutational data are the number of segregating sites, $C_{n}$ and the site frequency spectrum (SFS) $f_{n}:=(f(1),\dots,f(n-1))$, where $f(j)$ is the number of mutations detected in exactly $j$ sequences. We have $C_{n}=\sum_{j=1}^{n-1}f(j)$, and
[…]
The distribution of $C_{n}$ is known explicitly (Tavaré, 1984, eqn. (9.5)), but for the present purposes it is enough to note that $f(j)$ has approximately a Poisson distribution with mean $\theta/j$ (Dahmer and Kersting, 2015), and that $C_{n}$ has approximately a Poisson distribution with mean given in (2). The likelihood of observing $C_{n}=c$ given $\theta$ and $n$ is therefore approximately the probability that $\text{Poisson}\left(\theta\sum_{i=1}^{n-1}\frac{1}{i}\right)=c$. If the prior distribution for $\theta$ is $\pi(\theta)\sim$ Uniform$\left(a,b\right)$, the approximate posterior distribution of $\theta$ is
$\pi(\theta\mid c)\propto\mathbb{P}\left[\text{Poisson}\left(\theta\sum_{i=1}^{n-1}\frac{1}{i}\right)=c\right],\quad a<\theta<b.$ (3)
We compare the posterior distributions approximated with ABC-RF (package abcrf) and ABC-REJ (package EasyABC, (Jabot et al., 2023)) with uniform prior distribution U$\left(1,20\right)$ and summary statistics $S=C$. The observed data consists of one simulation with $\theta^{obs}\sim$ U$\left(1,20\right)$, resulting in $C^{obs}=34$ segregating sites. We configure ABC-REJ and ABC-RF to employ $N=$ 10,000 simulations each, from which ABC-REJ retains the 500 simulations with values closest to $C^{obs}$, and ABC-RF infers $\theta$ from $B=500$ trees. The posterior distributions from ABC-REJ (median = 4.65, mean = 4.82, variance = 1.46) and ABC-RF (median = 4.82, mean = 4.91, variance = 1.47) are close to the true distribution (Eq. 3; median = 4.58, mean = 4.68, variance = 0.67) (Figure 1a).
We next analyze the results when $S=\left\{C_{n},f(1),f(2),\dots,f(m)\right\}$ summarize the sequenced mutations; we chose $m=\lfloor\sqrt{n}\rfloor=31$. The individual $f(j)$ values contain significantly more noise and less information, especially for larger $j$. Implementation of ABC-REJ is therefore challenging, as the distance function $d(s^{*},s^{obs})$ must be calibrated to put more weight on $C_{n}$ and less on the $f(j)$, in order to avoid an extremely low acceptance rate. The lack of metrics in ABC-RF means that it does not face this issue, and its approximation for the posterior distribution (median = 4.29, mean = 4.31, variance = 0.89) is largely identical to the case where $S=C_{n}$ (Figure 1b). Moreover, abcrf’s variable importance analysis confirms that $C_{n}$ is the most important statistic in the inference, followed by the $f(j)$ in essentially increasing order of $j$ (Figure 1d).
We perform a final inference experiment using $S=\left\{f(1),f(2),\dots,f(m)\right\}$ as the observed statistics. Although ABC-RF’s inferred posterior distribution for $\theta$ has higher variance compared to the previous two cases, its center is still close to the true value (Figure 1c, median = 4.30, mean = 4.59, variance = 2.37).
These examples highlight the potential of random forests in the context of ABC inference. ABC-RF does not require calibrating the distance functions and tolerance thresholds, which have significant impact on the efficiency and accuracy in inferring the posterior distribution with other ABC methods. Furthermore, the robustness to noise in simulation statistics indicates reduced significance of selecting which statistics to summarize the observed data.
Figure 1: Inference of mutation rate in the coalescent model. a, b, c: Posterior distributions inferred with ABC-REJ and ABC-RF for mutation rate $\theta$, compared against the true posterior distribution (Eq. (3)). The statistics used in the inference consist of the mutation count $C$ (a), $C$ and the SFS $f(j)$ with $j=1,\dots,31$ (b) or SFS $f(j)$ with $j=1,\dots,31$ (c). d: ABC-RF’s variable importance analysis for summary statistics in the inference problem in b.
@@ 2.3 Distributional random forests
Choosing $(k_{*},s_{*})$ for multivariate problems then involves maximizing Eq. 4, aggregated across different parameters. Because this condition, denoted as CART criterion, only considers the difference in mean parameters in the child nodes and not the whole distributions, Ćevid et al. developed a second criterion based on the maximal mean discrepancy (MMD) metric (Ćevid et al., 2022). The MMD defines the difference between distributions of particles in two different sets by employing a positive-definite kernel and its embedding into a Reproducing Kernel Hilbert Space (RKHS) (Gretton et al., 2007). Due to the computational expense of MMD and the need to compute it for many different statistic and threshold candidates, the authors replace the kernel with its Fourier approximation. By default, they use the Gaussian kernel $k(\theta,\theta^{\prime})=\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{|\Theta|}}\exp\left(\frac{-\lVert\theta-\theta^{\prime}\rVert_{2}^{2}}{2\sigma^{2}}\right)$ with parameter count $|\Theta|$ and bandwidth $\sigma$, the median pairwise $L^{2}$ distance between the parameters of the particles in the parent node. With Fourier features $\omega_{1},\dots,\omega_{L}$ randomly chosen from the multivariate Gaussian distribution $\text{Normal}_{|\Theta|}\left(\textbf{0},\sigma^{-2}\textbf{I}_{|\Theta|}\right)$ where 0 and $\textbf{I}_{|\Theta|}$ are square zero and identity matrices of size $|\Theta|\times|\Theta|$, Ćevid et al. derive the approximate MMD as
@@
3       $R_{t}\leftarrow$ subsample of $R$ of size $n_{sample}$; $R_{t}$ is split into $R_{t1}$ and $R_{t2}$

### v2

@@ Abstract
Approximate Bayesian Computation (ABC) is a popular inference method when likelihoods are hard to come by. Practical bottlenecks of ABC applications include selecting statistics that summarize the data without losing too much information or introducing uncertainty, and choosing distance functions and tolerance thresholds that balance accuracy and computational efficiency. Recent studies have shown that ABC methods using random forest (RF) methodology perform well while circumventing many of ABC’s drawbacks. However, RF construction is computationally expensive for large numbers of trees and model simulations, and there can be high uncertainty in the posterior if the prior distribution is uninformative. Here we further adapt random forests to the ABC setting in two ways. The first exploits distributional random forests to provide a direct method for inferring the joint posterior distribution of parameters of interest, while the second describes a sequential Monte Carlo approach which updates the prior distribution iteratively to focus on the most likely regions in the parameter space. We show that the new methods can accurately infer posterior distributions for a wide range of deterministic and stochastic models in different scientific areas.
@@ 1 Introduction
Mathematical modeling has played an important role in studying scientific phenomena. Its practical applications often depend on accurately extracting model parameters $\theta$ from experimentally observed data, $y^{\rm obs}$. In the Bayesian framework, this entails inferring the posterior distribution
$\pi(\theta\mid y^{\rm obs})\propto f(y^{\rm obs}\mid\theta)\cdot\pi(\theta)$
@@
of the parameters from the data, exploiting the likelihood $f(y^{\rm obs}\mid\theta)$ and the prior $\pi(\theta)$ of such observations under the model with given parameter values. However, the likelihood function is often intractable to derive theoretically, difficult to compute numerically, or too complex to optimize directly. Approximate Bayesian Computation (ABC) was proposed as an alternative method to approximate the posterior distributions in such scenarios (Tavaré et al., 1997; Fu and Li, 1997; Pritchard et al., 1999; Beaumont et al., 2002). Using statistics to summarize model simulations, ABC seeks the parameters that result in minimal distance between the statistics and those of the observed data. Different variations have been proposed to improve the performance of the original ABC method, such as ABC Monte Carlo Markov Chain (Marjoram et al., 2003) and ABC Sequential Monte Carlo (Toni et al., 2008), and they have collectively found remarkable success in applications across different scientific areas (Sisson et al., 2018).
@@
The accuracy and efficiency of ABC in practice depends on several factors. First, one requires a distance function to compare observed and simulated data. The metric is typically weighted so that different statistics contribute equally (Prangle, 2017). However, an equally important consideration is the relevance of each statistic, defined as the amount of information that it carries toward identifying the underlying parameters. Optimizing these two criteria can be challenging (Jung and Marjoram, 2011). Second, ABC requires a tolerance threshold to decide whether proposed parameters are accepted or rejected. The tolerance $\epsilon$ poses a trade-off between the computational efficiency, which increases with higher $\epsilon$, and accuracy of the empirical posterior distributions, which improves as $\epsilon$ decreases. Choosing $\epsilon$ depends on the specific model, statistic choices and distance function, and it may require intuition or experimentation (Jung and Marjoram, 2011). Finally, and most importantly, ABC’s results depend on the choice of summary statistics. The approximated posterior distribution is only guaranteed to converge to the likelihood-based distribution as $\epsilon\to 0$ if the statistics are sufficient. For complex models, it is not always possible or realistic to find low-dimensional sufficient statistics (Sisson et al., 2018, Chapter 5). This can sometimes be remedied by inclusion of many distinct statistics. However, the posterior estimation may become distorted if the selected statistics are noisy or uninformative. These factors in combination mean that optimizing the performance of ABC in practice can be challenging and require extensive experimentation.
@@
We use some traditional ABC methods in comparisons with ABC-SMC-(D)RF described in the following sections. Recent ABC implementations using random forests are reviewed in Section 2, with illustrative examples in Section 3. In Section 4, we describe ABC-SMC-(D)RF’s methodology. We then demonstrate its performance, compared both to traditional ABC algorithms and previous RF-based methods, in a variety of deterministic and stochastic models (Section 5). We conclude with a discussion.
@@ 2 Random forests in the context of ABC
Although simple to implement, the ABC methods depend heavily on the choices of hyperparameters, including the summary statistic function $S(y)$, the distance function $d(s,s^{\prime})$ and the tolerance threshold $\epsilon$ (Sisson et al., 2018). Furthermore, the choices of Markov kernels to propose new particles in ABC Monte Carlo Markov Chain (ABC-MCMC) and ABC Sequential Monte Carlo (ABC-SMC) also affect the approximated posterior distributions. As random forests (RF) gain in popularity as a powerful non-parametric regression technique (Breiman, 2001; Segal, 2004; Rigatti, 2017; Desai and Ouarda, 2021), its applications in the context of ABC are also becoming more prominent. RF-based methods eliminate the need for choice of metric, tolerance level and the perturbation kernels, making them an attractive alternative to the previous ABC implementations. Importantly, the RF method is relatively robust to noise (Marin et al., 2018), in that $S(y)$ can include many poorly informative statistics without significant impact on the results.
@@ 2.1 One-dimensional ABC random forests
Raynal et al. introduced ABC random forest (ABC-RF) to approximate the posterior distribution for ABC inference problems with one parameter (Raynal et al., 2019), based on the random forest formulation (Algorithm 1) developed by Breiman (2001). The root node for each tree $T_{t}$ is the reference table, consisting of parameters drawn from the prior distribution and corresponding statistics. Each node $R_{*}$ in the decision tree is split into two nodes $R_{*1}$ and $R_{*2}$ by selecting one statistic $k_{*}\in\{1,\dots,|S|\}$, where $|S|$ is the total number of summary statistics, and a threshold $s_{*}$ such that the $L^{2}$-loss between the divided sets
@@
is minimized. Here $|R_{*}|$ is the number of particles in $R_{*}$, and $\bar{\theta}_{1}$, $\bar{\theta}_{2}$ are the parameter means in $R_{*1}$ and $R_{*2}$ respectively. The tree is finished when each leaf either consists of less than $N_{\rm min}$ particles, or all of the leaf’s particles have the same statistics. To guarantee forest diversity, the root node of each tree (Step 5 in Algorithm 1) is bootstrapped from the reference table, and the statistics considered for splitting each node (Step 9 in Algorithm 1) is limited to a randomly selected subset of size $n_{\rm try}$ among the available statistics $S$.
@@ 6  Tree $T_{t}$ is rooted in node $R_{0}$
7 while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{\rm min}$ do
8      Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{\rm min}$
9     Statistic candidates for splitting $R_{*}\leftarrow$ sample of size $n_{\rm try}$ among $S$
@@
Given an observation $y^{\rm obs}$, predicting its parameter $\theta$ based on tree $T_{t}$ involves following from the root node with the statistics $s^{\rm obs}$ and comparing $s^{\rm obs}_{k_{*}}$ with $s_{*}$ at each node $R_{*}$ to select the next node, until reaching a leaf $L_{t}(s^{\rm obs})$. The prediction for $\theta$ based on $T_{t}$ is then the average parameter among particles in $L_{t}(s^{\rm obs})$:
$\frac{\sum_{i=1}^{N}\theta^{(i)}\cdot n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}$
where $n_{t}^{(i)}$ is the number of times $\theta^{(i)}$ is duplicated in the bootstrapped sample $R_{0}$, and $\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}$ is the indicator for whether $s^{(i)}$ falls into the same leaf as $s^{\rm obs}$. The prediction for $\theta$ based on the whole forest is the average of predictions based on each tree:
$\sum_{i=1}^{N}\theta^{(i)}\cdot\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}$
Raynal et al. (2019) argued that this weighted estimate implies that the density of particles $\theta^{(i)}$ with corresponding weights $w^{(i)}$ (Algorithm 2) forms the approximation $\pi_{ABC-RF}\left(\theta\middle|s^{\rm obs}\right)$ for the posterior distribution.
@@
2   Follow the tree $T_{t}$ from root node with $s^{\rm obs}$ until locating leaf $L_{t}(s^{\rm obs})$
@@
4   Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{\rm obs})}}$
@@
Marin et al. (2022) developed an R package abcrf based on the ABC-RF method. For multivariate problems, ABC-RF is usually applied for each parameter marginally (Raynal et al., 2019). The authors also provided suggested values for RF hyperparameters; by default, $N_{\rm min}=5$ and $n_{\rm try}=|S|/3$. Among the most important hyperparameters is the count of trees $B$, for which they recommend analyzing whether the out-of-bag mean squared error stabilizes around the selected $B$ (Raynal et al., 2019; Pudlo et al., 2016).
@@ 2.2 Distributional random forests
Choosing $(k_{*},s_{*})$ for multivariate problems then involves maximizing Eq. 2, aggregated across different parameters. Because this condition, denoted as the CART criterion, only considers the difference in mean parameters in the child nodes and not the whole distributions, Ćevid et al. developed a second criterion based on the maximal mean discrepancy (MMD) metric (Ćevid et al., 2022). The MMD defines the difference between distributions of particles in two different sets by employing a positive-definite kernel and its embedding into a Reproducing Kernel Hilbert Space (RKHS) (Gretton et al., 2007). Due to the computational expense of MMD and the need to compute it for many different statistic and threshold candidates, the authors replace the kernel with its Fourier approximation. By default, they use the Gaussian kernel $k(\theta,\theta^{\prime})=\left.\exp\left(-\lVert\theta-\theta^{\prime}\rVert_{2}^{2}\middle/2\sigma^{2}\right)\middle/\left(\sigma\sqrt{2\pi}\right)^{|\Theta|}\right$. with parameter count $|\Theta|$ and bandwidth $\sigma$, the median pairwise $L^{2}$ distance between the parameters of the particles in the parent node. With Fourier features $\omega_{1},\dots,\omega_{L}$ randomly chosen from the multivariate Gaussian distribution $\text{Normal}_{|\Theta|}\left(\textbf{0},\sigma^{-2}\textbf{I}_{|\Theta|}\right)$ where 0 and $\textbf{I}_{|\Theta|}$ are square zero and identity matrices of size $|\Theta|\times|\Theta|$, Ćevid et al. derive the approximate MMD as
@@
3 $R_{t}\leftarrow$ subsample of $R$ of size $n_{\rm sample}$; $R_{t}$ is split into $R_{t1}$ and $R_{t2}$